how long does it take a quarter dropped from the roof of a building 224 feet high to reach the group, (if there were no air resistance by using the formula
t=1/4 square d where d is the distance in feet and t is time in seconds

Answer by Alan3354(69443)

how long does it take a quarter dropped from the roof of a building 224 feet high to reach the group, (if there were no air resistance by using the formula
t=1/4 square d where d is the distance in feet and t is time in seconds
------------------
t = (1/4)^2*d
t = 1/16*224 = 14 seconds
-----------
What sort of group?